OpenStudy (anonymous):

66 2/3 of what is 159 ?

OpenStudy (anonymous):

0.4192872075471698

OpenStudy (stacey):

To solve a problem like this change it to an algebra problem. "fraction of something" means that the fraction is multiplied times the something. "is" is the equal sign. 66 2/3 * x = 159. x=159 / (66 2/3) = 2.385